Translate the following verbal phrases into mathematical phrases
1. Twice the quotient of 4 times a number and 8
2. Fourteen plus the product of 3 less than a number and 10
3. A number plus the product of 7 and the difference of the number and 5
4. Three less than the sum of a number and 10
5. The square of z
6. The cube of m Answer by math_tutor2020(3838) 

In the future, please post one problem at a time.
I'll do the first three questions to get you started.
-----------------
Problem 1
x = some unknown number
4x = four times that number
4x/8 = quotient of 4x and 8
2*(4x/8) = twice that previous amount
This can be simplified, but I'll leave it un-simplified so it fits the question format.
-----------------
Problem 2
x = some unknown number
x-3 = three less than the number
10(x-3) = product of 10 and (x-3)
14+10(x-3) = add the previous expression to 14
As with the previous answer, I'll leave it un-simplified.
-----------------
Problem 3
x = some unknown number
x-5 = difference of the number and 5
7(x-5) = product of 7 and the previous amount
x+7(x-5) = add the previous expression to x
As with the previous two answers, I'll leave it un-simplified.